OpenStudy (anonymous):

How can I memorize square roots 1-30? I got the list from yahoo answers: 1² = 1 2² = 4 3² = 9 4² = 16 5² = 25 6² = 36 7² = 49 8² = 64 9² = 81 10² = 100 11² = 121 12² = 144 13² = 169 14² = 196 15² = 225 16² = 256 17² = 289 18² = 324 19² = 361 20² = 400 21² = 441 22² = 484 23² = 529 24² = 576 25² = 625 26² = 676 27² = 729 28² = 784 29² = 841 30² = 900 but any ideas/ways on how i could memorize them? I need to know them by monday.

OpenStudy (agent47):

Look at the bright side, you don't have to memorize 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 20, 30

OpenStudy (agent47):

And I only know 1-13, 15-17, 20, 25, 30 anyway, I'm not sure why you'd want to know all of them. The ones I know got me through my math degree.

OpenStudy (anonymous):

Agent is correct, memorize the right side of the equal sign, in that order. I guess it might help. Methods of memorizing vary through the student body. You have to find the method that's suitable for you.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

You can look for patterns. Also, knowing how to FOIL binomials helps.

OpenStudy (anonymous):

NP, what do you mean?

OpenStudy (anonymous):

Did you notice that all those square numbers differ from one-another by an odd number?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

Look at the first few squares: 0, 1, 4, 9, 16, 25 The difference between 0 and 1 is 1 The difference between 1 and 4 is 3 The difference between 4 and 9 is 5 and so on.

OpenStudy (anonymous):

i see.

OpenStudy (anonymous):

The pattern goes: +1, +3, +5, +7, +9, +11, +13, + ... Always increasing by an odd number.

OpenStudy (anonymous):

So let's say you forget what 13^2 is, but you remember that 12^2 = 144...

OpenStudy (anonymous):

ah. okay. thanks for the tip.

OpenStudy (anonymous):

that is a good way of trying to memorize them.

OpenStudy (anonymous):

To find 13^2 from 12^2, find the 13th odd number (2*13-1 = 25) 144 + 25 = 169. ;-)

OpenStudy (anonymous):

mhmmm, could you quiz me in a bit?

OpenStudy (anonymous):

also, is there another pattern?

OpenStudy (anonymous):

What I meant by FOILing binomials (this is something you learn in algebra, but can be used for arithmetic as well) is FOIL = First Outer Inner Last So for, say, 16^2 --> think of it as (10+6)*(10+6) First: 10*10 = 100 Outer: 10*6 = 60 Inner: 6*10 = 60 (same as the outer if you're squaring so just double up what you just got. Last: 6*6 = 36 Add it all up: 100+120+36 = 256. With practice, you can do this quickly in your head. You really only need to memorize the squares up to 10^2 (but you should memorize at least up to 20 or 25 if you can).

OpenStudy (anonymous):

im in pre algebra

OpenStudy (anonymous):

but thanks.

OpenStudy (anonymous):

Then you at least know order of operations and distributive property - that's all that is.

OpenStudy (anonymous):

I recommend making flash cards and practice mental arithmetic whenever you can. Resist the temptation to reach for a calculator whenever you have some arithmetic to do. Try it mentally first - if it takes you more than a few seconds and you start to forget your last calculation, then write it down. Only if that fails should you use a calculator.

OpenStudy (anonymous):

What I like to do for practice is to take a deck of cards - you can take out the face cards and aces, so you just have 2-10 - shuffle them and deal out two numbers and try to add them, subtract them, multiply them, and divide them as quickly as you can mentally, and then deal out two more . . . until you go through the whole deck.

OpenStudy (anonymous):

Cliffs, ill be back in about 30 minutes. do you think you will still he online?

OpenStudy (anonymous):

be*

OpenStudy (anonymous):

Maybe. I'm about to get ready for bed, but if I'm off elsewhere, try tagging my name "@CliffSedge"

OpenStudy (anonymous):

Ok,thanks.

OpenStudy (anonymous):

see you cliffs.

OpenStudy (anonymous):

ill log out for right now. then log back in when i am back.

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

What's up?

OpenStudy (anonymous):

nothing much.

OpenStudy (anonymous):

you?

OpenStudy (anonymous):

About to go to bed. Pop quiz: (mentally, no pencil, no calculator) find 40^2

OpenStudy (anonymous):

160?

OpenStudy (anonymous):

On the right track, but not quite.

OpenStudy (anonymous):

4x4 16 add a 0

OpenStudy (anonymous):

40=4*10 4*4=16 10*10=100

OpenStudy (anonymous):

1600

OpenStudy (anonymous):

Right, now look at 20^2 and 30^2 and see that they are done the same way.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

400

OpenStudy (anonymous):

This is a key strategy to be able to figure things out mentally an to help you memorize patterns. Next example: Find 35^2, knowing that 35=30+5 (Remember, FOIL)

OpenStudy (anonymous):

35x35

OpenStudy (anonymous):

1825

OpenStudy (anonymous):

Seems a little too big. How did you find that?

OpenStudy (anonymous):

see, i do not understamd foil, because i haventnbeen taught that yet, so i attempted to do some quick multiplication.

OpenStudy (anonymous):

That's fine. It just takes practice. 35*35 = (30+5)*(30+5) Using distributive property: = 30*30 + 30*5 + 30*5 +5*5 30*30 = 900 5*5=25 The 30*5 shows up twice, so just double that to 300, and the total is 900+300+25.

OpenStudy (anonymous):

You'll only have to use this trick if you trying to mentally figure out squares which you've forgotten. For example, I can remember most squares up to 25, but sometimes I forget some, like 19^2 or 23^2, so I do FOIL in my head. You should definitely strive to memorize all squares up to 16^2; get those down solid first - they're pretty easy. After that, work on 17^2 through 25^2. Try the FOIL trick and the pattern of adding odd numbers along the way. Once you have those, just practice once in a while, either with flash cards, or try to just name all the squares you can in order and see how high you can go.

OpenStudy (anonymous):

cliffs, you'll have to excuse me - gotta go. thanks for helping out :)

OpenStudy (anonymous):

Sure thing, me too. Have a good night, Rachel.

OpenStudy (anonymous):

you too, cliffs.